# On the use of quantality in nuclei and many-body systems

### v1

@@ VI Conclusions
When related to the dimensionless coupling constant, the quantality drives the spin-orbit effect in finite systems. The nucleus is a very specific system, due to its naturalness (i.e. its dimensionless constants of the order of the unity). The nuclear clusterization condition could even be considered as a supernaturalness state: all the dimensionless parameters are close to one, including the localization parameter. The impact of quantality on the spin-orbit effect is exhibited, using the spin-orbit rule obtained from a non-relativistic reduction of the Dirac equation. Also, the quantality is shown to drive the binding energy per particle compared to the shell gap.
Three additional dimensionless quantities have been discussed. It should be reminded that the quantities used in the present work, contain information from the solved Schrödinger equation (of course, in an approximated way), through e.g. the evaluation of the dispersion of the nucleonic wave function obtained by microscopic EDF or HO approximation, or the behavior of the spin-orbit effect through a non-relativistic reduction of the Dirac equation. The present results show how quantality can be useful to nuclear structure, beyond a mere evaluation of its value in various many-body systems.

### v2

@@ VI Conclusions
When related to the dimensionless coupling constant, the quantality drives the spin-orbit effect in finite systems. The nucleus is a very specific system: its dimensionless constants are close to the order of the unity. The nuclear clusterization condition could even be considered as a superspecificness state: all the dimensionless parameters are close to one, including the localization parameter. The impact of quantality on the spin-orbit effect is exhibited, using the spin-orbit rule obtained from a non-relativistic reduction of the Dirac equation. Also, the quantality is shown to drive the binding energy per particle compared to the shell gap.
Three additional dimensionless quantities have been discussed. It should be reminded that the quantities used in the present work, contain information from the solved Schrödinger equation (of course, in an approximated way), through e.g. the evaluation of the dispersion of the nucleonic wave function obtained by microscopic EDF or HO approximation, or the behavior of the spin-orbit effect through a non-relativistic reduction of the Dirac equation. The present results show how quantality can be useful to nuclear structure, beyond a mere evaluation of its value in various many-body systems. It can be evaluated in microscopic approaches such as EDF or Ab initio and EFT ones.